*Frozen (quarter-pound) hamburger patties cost $5.80 for 12. How many can be purchased for $7.00?
If you could show your work that would really help.
Thanks so much

Answer by edjones(8007)

x/7=12/5.8 Proportion
x=84/5.8 multiply each side by 7
=14.48... patties
Did you make a mistake copying the problem?
.
Ed